# Exact low-dimensional description for fast neural oscillations with low-firing rates

### 14 Dec 2023 v3

@@ III Exact Firing Rate Model
In the thermodynamic limit, Eqs. (1,3) with $\Gamma=0$ are exactly described by a low-dimensional system of differential equations that we refer to as QIF-FRM Montbrió et al. (2015). Inspired by recent results in the context of the Kuramoto model Tönjes and Pikovsky (2020); Tanaka (2020), we next show that an identical set of exact FRM is obtained if, in addition to heterogeneity, neurons are driven by independent Cauchy noise  [^4] See also recent attempts to obtain approximated mean field descriptions of populations of QIF neurons with independent Gaussian noise  Goldobin et al. (2021); Ratas and Pyragas (2019); Goldobin (2021); di Volo et al. (2022)..
@@ III.2 Mean membrane potential and mean firing rate
Finally, replacing $y=v$, $x=r$, and $\eta=\overline{\eta}-\Delta i$ in Eq (15) and taking real and imaginary part leads to
@@ Acknowledgements.
The authors thank Jordi Garcia-Ojalvo for helpful discussions. PC acknowledges financial support from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 101017716 (Neurotwin). EM acknowledges support by the Agencia Estatal de Investigación under the Project No. PID2019-109918GB-I00. Note added.—During the review process of this article, other Authors have published the article Ref. Pietras et al. (2023), which introduces a low-dimensional reduction for ensembles of theta neurons beyond the Lorentzian ansatz used here.
@@ B.2 Numerical results with Gaussian heterogeneity and noise
In figure 6(a) we show a bifurcation diagram for QIF neurons with $q$-Gaussian heterogeneites for different values of $n$. We use the half-width at half-maximum (HWHM) $d$ as control parameter of the heterogeneity. Notice that $d=\Delta$ for a Cauchy distribution ($n=1$) and $d=\sigma\sqrt{2\log(2)}$ for a Gaussian distribution with standard deviation $\sigma$ ($n\to\infty$). Black line in figure 6 corresponds to the black continuous line in Fig. 3. As $n$ increases, the region of oscillations widens. The colormap shows the level of activity in the region of oscillations obtained using simulations of Eqs. (1,3), but considering $\eta_{j}$ to be distributed as a Gaussian with mean $\overline{\eta}$ and standard deviation $\sigma$. In spite of the enlargement of the instability region, the Hopf bifurcation remains displaying two different cases: For low coupling $J$ most neurons remain active even at the bifurcation line ($\Omega\approx\langle r\rangle$). Instead, for high coupling there is a large degree of suppression as $d$ increases.
For Gaussian noise, attempts to derive mean-field equations have been put forward Goldobin et al. (2021); Ratas and Pyragas (2019); Goldobin (2021); di Volo et al. (2022). However, these theories build on weak noise approximations ($\sigma\ll 1$) and are thus unsuitable to analyze networks with large fluctuations. Figure 6(b) displays a numerical bifurcation diagram of the QIF network (Eqs. (1,3)) with $\xi_{i}(t)$ being Gaussian white noise with standard deviation $\sigma=d/\sqrt{2\log(2)}$. The scenario remains remarkably similar to the case of Gaussian heterogeneity, with still two qualitatively different transitions towards stationarity as $d$ is increased. It is worth noting that we did not find cluster states. This contrasts with the results of Brunel and Hansel (2006), which analyze a similar setup with other integrate-and-fire models and find cluster instabilities for very low levels of Gaussian noise. A possible explanation might be the lack of a fixed delay and/or a rise synaptic time in our modelling setup.
Figure 6: Phase diagram of the QIF network with Gaussian distributions. (a) Lines: Supercritical Hopf bifurcations of the mean-field equations in Pyragas and Pyragas (2022) for different values of $n$ (parameter values as in our paper). The colormap corresponds to the average number of spikes per oscillation cycle for a QIF network with Gaussian heterogeneity. It has been numerically obtained through numerical integration of the QIF network Eqs. (22,3) of the main paper ($N=8192$). (b) Black line as in Fig. 3 of the main paper. Colormap corresponds to simulations of QIF neurons ($N=8192$) with Gaussian white noise and without heterogeneity (rest of the parameters as in the body of the paper). In both panels, a simulation has been considered to be in the oscillatory regime if the standard deviation of the synaptic activity $s(t)$ is below $10^{-3}$.

### 01 Feb 2024 v4

# Exact low-dimensional description for fast neural oscillations with low firing rates

@@ III Exact Firing Rate Model
In the thermodynamic limit, Eqs. (1,3) with $\Gamma=0$ are exactly described by a low-dimensional system of differential equations that we refer to as QIF-FRM Montbrió et al. (2015). Inspired by recent results in the context of the Kuramoto model Tönjes and Pikovsky (2020); Tanaka (2020), we next show that an identical set of exact FRM is obtained if, in addition to heterogeneity, neurons are driven by independent Cauchy noise 444 See Ref. Pietras et al. (2023) for an alternative derivation. Additionally, see also recent attempts to obtain approximated mean field descriptions of populations of QIF neurons with independent Gaussian noise  Goldobin et al. (2021); Ratas and Pyragas (2019); Goldobin (2021); di Volo et al. (2022)..
@@ III.2 Mean membrane potential and mean firing rate
Finally, replacing $y=v$, $x=\pi\tau_{m}r$, and $\eta=\overline{\eta}-\Delta i$ in Eq (15) and taking real and imaginary part leads to
@@ Acknowledgements.
The authors thank Jordi Garcia-Ojalvo for helpful discussions. PC acknowledges financial support from the European Union’s Horizon 2020 research and innovation programme under grant agreement No 101017716 (Neurotwin). EM acknowledges support by the Agencia Estatal de Investigación under the Project No. PID2019-109918GB-I00.
@@ B.2 Numerical results with Gaussian heterogeneity and noise
In figure 6(a) we show a bifurcation diagram for QIF neurons with $q$-Gaussian heterogeneites for different values of $n$. We use the half-width at half-maximum (HWHM) $d$ as control parameter of the heterogeneity. Notice that $d=\Delta$ for a Cauchy distribution ($n=1$) and $d=\sigma\sqrt{2\ln(2)}$ for a Gaussian distribution with standard deviation $\sigma$ ($n\to\infty$). Black line in figure 6 corresponds to the black continuous line in Fig. 3. As $n$ increases, the region of oscillations widens. The colormap shows the level of activity in the region of oscillations obtained using simulations of Eqs. (1,3), but considering $\eta_{j}$ to be distributed as a Gaussian with mean $\overline{\eta}$ and standard deviation $\sigma$. In spite of the enlargement of the instability region, the Hopf bifurcation remains displaying two different cases: For low coupling $J$ most neurons remain active even at the bifurcation line ($\Omega\approx\langle r\rangle$). Instead, for high coupling there is a large degree of suppression as $d$ increases.
For Gaussian noise, attempts to derive mean-field equations have been put forward Goldobin et al. (2021); Ratas and Pyragas (2019); Goldobin (2021); di Volo et al. (2022). However, these theories build on weak noise approximations ($\sigma\ll 1$) and are thus unsuitable to analyze networks with large fluctuations. Figure 6(b) displays a numerical bifurcation diagram of the QIF network (Eqs. (1,3)) with $\xi_{i}(t)$ being Gaussian white noise with standard deviation $\sigma=d/\sqrt{2\ln(2)}$. The scenario remains remarkably similar to the case of Gaussian heterogeneity, with still two qualitatively different transitions towards stationarity as $d$ is increased. It is worth noting that we did not find cluster states. This contrasts with the results of Brunel and Hansel (2006), which analyze a similar setup with other integrate-and-fire models and find cluster instabilities for very low levels of Gaussian noise. A possible explanation might be the lack of a fixed delay and/or a rise synaptic time in our modelling setup.
Figure 6: Phase diagram of the QIF network with Gaussian distributions. (a) Lines: Supercritical Hopf bifurcations of the mean-field equations in Pyragas and Pyragas (2022) for different values of $n$ (parameter values as in Fig. 3). The colormap corresponds to the average number of spikes per oscillation cycle for a QIF network with Gaussian heterogeneity. It has been numerically obtained through numerical integration of the QIF network Eqs. (22,3) of the main paper ($N=8192$). (b) Black line as in Fig. 3 of the main paper. Colormap corresponds to simulations of QIF neurons ($N=8192$) with Gaussian white noise and without heterogeneity (rest of the parameters as in the body of the paper). In both panels, a simulation has been considered to be in the oscillatory regime if the standard deviation of the synaptic activity $s(t)$ is below $10^{-3}$.